# Simulating thermodynamic properties of dinuclear metal complexes using Variational Quantum Algorithms

### v1

@@ V Experimental Comparison: Thermodynamic properties of $NH_{4}CuPO_{4}\cdot H_{2}O$
In order to validate our simulation, we conduct a study of the $NH_{4}CuPO_{4}\cdot H_{2}O$ compound Koo and Whangbo (2008); Chakraborty et al. (2014); Nawa et al. (2012); Pujana et al. (1998), a dinuclear Cu(II) metal complex using the proposed variational quantum algorithm. This material appears as a model in scientific literature to study quantum correlations and behaviors in condensed matter systems Chakraborty et al. (2014). In this regard, the molar magnetic susceptibility was investigated in the temperature range of 2 K to 100 K, showing antiferromagnetic correlations in the system Koo and Whangbo (2008); Nawa et al. (2012); Pujana et al. (1998). The experimental data were analyzed using an isolated Heisenberg dimer model, allowing the determination of the exchange coupling constant $J$ Pujana et al. (1998). In addition, specific heat measurements were conducted throughout a temperature range of 2 K to 10 K, yielding valuable information about the thermodynamic properties of the material at low temperatures Pujana et al. (1998); Chakraborty et al. (2014).
@@
Figure 7: (a) Magnetic susceptibility and (b) Magnetic specific heat for $NH_{4}CuPO_{4}\cdot H_{2}O$ as a function of temperature. The filled red points were obtained using the Variational Quantum Thermalizer Algorithm (VQT) for $J_{VQA}/k_{B}=4.18\leavevmode\nobreak\ K$, compared to the experimental data (unfilled blue points) $J_{exp}/k_{B}=5\leavevmode\nobreak\ K$ extracted from the reference Chakraborty et al. (2014); Pujana et al. (1998). The solid lines represent the theoretical models described in Eqs. (16) and (19).
@@
As can be seen, the proposed variational quantum algorithm accurately reproduces the experimental data for molar magnetic susceptibility and specific heat. The value of the magnetic coupling parameter, obtained from the variational quantum algorithm, that best fit the experimental data was $J_{VQA}/k_{B}=4.18\leavevmode\nobreak\ K$. This result is in good agreement with the fit of the experimental data $J_{exp}/k_{B}=5\leavevmode\nobreak\ K$ reported in the literature Chakraborty et al. (2014); Pujana et al. (1998). Additionally, the specific heat data obtained from the simulation showed a peak of the Schottky anomaly (as described in sections III) at around $3.5\leavevmode\nobreak\ K$, which is consistent with the reported experimental data Chakraborty et al. (2014). The close agreement between the simulated and experimental results indicates the potential of quantum algorithms in material science research. This suggests that the quantum algorithm used in the simulation is effective in predicting the behavior of magnetic materials. Further studies can explore other systems to validate the accuracy and reliability of quantum algorithms in material science.
@@ VI Conclusions
In summary, this work showcases the significant potential of Variational Quantum Algorithms (VQAs) in simulating the thermodynamic properties of advanced materials, particularly focusing on dinuclear spin $1/2$ metal complexes. Using an adapted version of the Variational Quantum Thermalizer algorithm, we simulated the evolution of thermal properties with changing temperatures, highlighting the transition from quantum to classical behavior as temperature increases. The simulated results provide valuable insights into the thermodynamic properties of the Heisenberg dimer system, offering a deeper understanding of its behavior at the quantum level. Additionally, we compare the prediction of the molar magnetic susceptibility and the specific heat obtained from the proposed VQA with experimental data reported in the literature for the $NH_{4}CuPO_{4}\cdot H_{2}O$ dinuclear metal complex. The agreement between the simulated and experimental results validates the accuracy of the VQA method in predicting material properties, demonstrating the promise of quantum algorithms for advancing research in materials science. Therefore, this approach demonstrates the potential of quantum computing in accurately predicting and analyzing material properties, paving the way for more efficient and innovative material design processes.
@@ Acknowledgements.
The authors thank M. S. Reis for his helpful comments and the Bahia State Research Support Foundation (FAPESB) for financial support.

### v2

@@ V Experimental Comparison: Thermodynamic properties of $NH_{4}CuPO_{4}\cdot H_{2}O$
In order to validate our simulation, we conduct a study of the $NH_{4}CuPO_{4}\cdot H_{2}O$ compound Koo and Whangbo (2008); Chakraborty et al. (2014); Nawa et al. (2012); Pujana et al. (1998), a dinuclear Cu(II) metal complex using the proposed variational quantum algorithm. This material appears as a model in scientific literature to study quantum correlations and behaviors in condensed matter systems Chakraborty et al. (2014). In this regard, the molar magnetic susceptibility was investigated in the temperature range of 2 K to 100 K, showing antiferromagnetic correlations in the system Koo and Whangbo (2008); Nawa et al. (2012); Pujana et al. (1998). The experimental data were analyzed using an isolated Heisenberg dimer model, allowing the determination of the exchange coupling constant $J$ Pujana et al. (1998). In addition, specific heat measurements were conducted throughout a temperature range of 2 K to 100 K, yielding valuable information about the thermodynamic properties of the material at low temperatures Pujana et al. (1998); Chakraborty et al. (2014).
@@
Figure 6: (a) Magnetic susceptibility and (b) Magnetic specific heat for $NH_{4}CuPO_{4}\cdot H_{2}O$ as a function of temperature. The unfilled red points were obtained using the Variational Quantum Thermalizer Algorithm (VQT) for $J_{VQA}/k_{B}=4.2\leavevmode\nobreak\ K$, compared to the experimental data (filled blue points) $J_{exp}/k_{B}=5\leavevmode\nobreak\ K$. The dashed line represents the theoretical models described in Eqs. (12) and (15).
@@
As can be seen, the proposed variational quantum algorithm accurately reproduces the experimental data for molar magnetic susceptibility and specific heat. The value of the magnetic coupling parameter, obtained from the variational quantum algorithm, that best fit the experimental data was $J_{VQA}/k_{B}=4.2\leavevmode\nobreak\ K$. This result is in good agreement with the fit of the experimental data $J_{exp}/k_{B}=5\leavevmode\nobreak\ K$ reported in the literature Chakraborty et al. (2014); Pujana et al. (1998). Additionally, the specific heat data obtained from the simulation showed a peak of the Schottky anomaly (as described in sections III) at around $3.5\leavevmode\nobreak\ K$, which is consistent with the reported experimental data Chakraborty et al. (2014). The close agreement between the simulated and experimental results indicates the potential of quantum algorithms in material science research. This suggests that the quantum algorithm used in the simulation is effective in predicting the behavior of magnetic materials.
However, it is worth noting that the results, even when simulated in an ideal environment, show slight discrepancies with theoretical and experimental results. In this regard, it is important to highlight some of the inherent limitations of variational algorithms: the gradient function in classical optimization aims to find the better parameters $\phi$ that minimize the cost function $C$ (eq. 2), but the convergence of these parameters is not always exact. Some effects must be taken into account, such as barren plateaus noise Larocca et al. (2024). This effect refers to flat regions (or plateaus) in the cost function optimization landscape of these algorithms, where the gradient of the cost function is very small or almost zero Wang et al. (2021); Uvarov and Biamonte (2021). The deeper and larger the circuit, the more likely these noises are Kulshrestha and Safro (2022). On the other hand, some alternatives to accelerate variational quantum eigensolvers involve integrating advanced Bell measurement techniques Cao et al. (2024). These methods aim to improve efficiency, thereby reducing computational time, which can facilitate advancements in quantum chemistry and material science.
V.1 Realistic noisy simulation
Conversely, in order to analyze the algorithm viability in current quantum devices, the VQT was simulated in shot-based and noisy simulators. For this purpose, weused the IBM Quantum Experience simulators, qasm_simulator and FakeSheerBrooke. The qasm_simulator in Qiskit is a classical backend for simulating quantum circuits in a shot-based environment described in QASM format Javadi-Abhari et al. (2024). FakeSheerBrooke is a noisy simulator in Qiskit that mimics the noise characteristics of 127-qubits IBM’s Sherbrooke real quantum computer.
To obtain the expected values to be minimized by the optimization process, it is necessary to measure the observable at their respective bases in the quantum circuit. It can be done by adding rotation gates after calculating the state in the ansatz Koch et al. (2020). Naturally, qiskit measures values in the computational base, corresponding to the Z base. For the base X, it is necessary to add Hadamard gates on both qubits, while for the base y it is necessary to add a rotation on the x axis with an angle $-\pi/2$. In each circuit we choose 1024 shots to the circuits to measure the expected value of each observable in the system Hamiltonian (8).
In each circuit corresponding to an observable of the Heisenberg Hamiltonian, we choose 8192 shots, which is the default of IBM Qiskit itself. For the process of optimizing the ansatz parameters, the same minimization method was used, but adding a limit of 400 iterations at most, taking into account the longer time needed to compile the algorithm in noisy simulators. To arrive at these configurations, tests were carried out with different numbers of shots and the maximum number of iterations. Figure 7 shows the best test results obtained with the previously specified settings in a single run.
Figure 7: (a) Magnetic susceptibility and (b) Magnetic specific heat for $NH_{4}CuPO_{4}\cdot H_{2}O$ as a function of temperature obtained in a single run of the algorithm in a noisy simulator. The curves show the simulation performance for the shot-based noisy model (orange x-shaped data) in relation to the simulation run in the shot-based qasm_simulator without noisy (filled blue points) and in the StateVector simulator (unfilled red points) for $J_{VQA}/k_{B}=4.2\leavevmode\nobreak\ K$. The dashed line represents the theoretical models described in Eqs. (12) and (15). In each circuit, we choose 8192 shots and 400 simulation steps in classical optimization.
The results show the impact of shot-based and noisy in algorithm in our simulation. When using the qasm_simulator, it is possible to observes small discrepancies compared to the Statevector simulation. Even without noise, this outcome is expected because the shot-based process generates a probability distribution for each eigenstate based on the number of measurements taken Li et al. (2023). The more measurements taken (shots), the more accurate the probability distribution becomes. In our algorithm, we used 8192 shots, which provided a reasonable approximation of the state vector.
As expected, the FakeSherbrooke simulator (Noise Model) demonstrated greater discrepancies compared to others. This is a direct result of the fact that the simulator incorporates inherent characteristics of the real Sherbrooke quantum computer. Among these characteristics, those with the greatest influence on the result are: influence of disconnection between qubits Kapit et al. (2020), decoherence related to the rate of loss of quantum information Shaji and Caves (2007), gate error due to a possible failure in its implementation Ayral et al. (2021) or noise inherent to measurements in the circuit Resch and Karpuzcu (2021). However, even with these noise effects, it is possible to notice that the outcomes follow the behavior of the theoretical model, with the points being concentrated in regions close to the theoretical curve.
In summary, the use of quantum simulators is crucial for the progression of quantum computing, as they enable researchers to foresee and address practical issues prior to executing algorithms on actual hardware . Simulators play a significant role in understanding and mitigating the effects of noise, thereby aiding in the development of more robust and accurate quantum computers. The findings from our study highlight the discrepancies and challenges associated with shot-based and noisy simulations, underscoring the importance of these tools in the pursuit of fault-tolerant quantum computing. As the field advances, ongoing refinement of simulation techniques and noise models will be essential for bridging the gap between theoretical models and practical implementations, particularly in the study of thermodynamic properties of advanced materials using quantum computing.
@@ VI Conclusions
In summary, this work showcases the significant potential of Variational Quantum Algorithms (VQAs) in simulating the thermodynamic properties of advanced materials, particularly focusing on dinuclear spin $1/2$ metal complexes. Using the Variational Quantum Thermalizer algorithm, we simulated the evolution of thermal properties with changing temperatures, highlighting the transition from quantum to classical behavior as temperature increases. The simulated results provide valuable insights into the thermodynamic properties of the Heisenberg dimer system, offering a deeper understanding of its behavior at the quantum level. Additionally, we compare the prediction of the molar magnetic susceptibility and the specific heat obtained from the proposed VQA with experimental data reported in the literature for the $NH_{4}CuPO_{4}\cdot H_{2}O$ dinuclear metal complex. The agreement between the simulated and experimental results validates the accuracy of the VQA method in predicting material properties, demonstrating the promise of quantum algorithms for advancing research in materials science. Therefore, this approach demonstrates the potential of quantum computing in accurately predicting and analyzing material properties, paving the way for more efficient and innovative material design processes.
@@ Acknowledgements.
The authors thank M. S. Reis for his helpful comments and Fundação de Amparo à Pesquisa do Estado da Bahia - FAPESB for its financial support (grant number APP0041/2023). Lucas Q. Galvão thanks EMBRAPII and the Brazilian Ministry for Science, Technology and Innovation (MCTI) for financial support.
Appendix A HAMILTONIAN AND DENSITY MATRICES TOMOGRAPHY
Figure 8 depicts the tomography of the system Hamiltonian in the computational basis, as defined by Eq. (7). As can be seen, the matrix entries are parameterized by the system’s magnetic coupling constant $J$.
Figure 8: Simulated Hamiltonian matrix, parametrized by the magnetic coupling constant $J$, in the computational basis.
Figure 9: Thermal states obtained in the simulation for different temperature values: (a) thermal state at $T=5\times 10^{-4}\leavevmode\nobreak\ K$; (b) $T=10\leavevmode\nobreak\ K$; (c) $T=20\leavevmode\nobreak\ K$; (d) $T=300\leavevmode\nobreak\ K$. All these thermal states were reproduced after $400$ simulation steps, with $J/k_{B}=1\leavevmode\nobreak\ K$, where $J$ and $k_{B}$ are the magnetic coupling and the Boltzmann constant, respectively.
As can be seen, considering $J/k_{B}=1\leavevmode\nobreak\ K$, Figure 9 (a) shows that at very low temperatures ($T=5\times 10^{-4}\leavevmode\nobreak\ K$), the system exists in the pure singlet ground state $|\varrho_{4}\rangle=\frac{1}{\sqrt{2}}\left(|01\rangle{-}|10\rangle\right)$, which is a maximally entangled state. As we increase the temperature to $10\leavevmode\nobreak\ K$ in Figure 9 (b), the density matrix approaches a two-qubit X-shaped state Sarandy (2009), which is characterized by the terms of the secondary diagonal of the matrix. Upon reaching $20\leavevmode\nobreak\ K$ in Figure 9 (c), the populations increase while the coherence terms decrease. Finally, when the temperature reaches a high enough value of $T=300\leavevmode\nobreak\ K$, the system tends towards a classical probability distribution, which is characterized by the density matrix in its diagonal form, as shown in Figure 9 (d). This evolution from an entangled mixed state to a classical probability distribution highlights the transition from quantum to classical behavior as temperature increases. The changes in the density matrix elements reflect the diminishing quantum correlations and the emergence of classical-like features in the system as thermal fluctuations dominate Schlosshauer (2019).